# Quantization without Tears

### v3

@@ 4 Experiments
Percentile[23] 4/4 014.0 68.4
Percentile + QwT 4/4 016.0 74.5
Percentile + QwT 4/4 016.0 75.8
@@
Percentile[23] 6/6 019.9 76.0
Percentile + QwT 6/6 021.9 76.8
Percentile + QwT 6/6 021.9 76.8
@@ 4.1 Experiments on Image Classification
Results on different backbones. Table 1 summarizes the quantization results when applying QwT across different backbone architectures. Specifically, we selected RepQ-ViT [27] and Percentile [23] as the baseline methods for the Transformer family [11, 49, 34] and ResNet [17], respectively. The results show that incorporating the QwT module consistently boosts the recognition accuracy, leading to an average increase of approximately 2.6%, and even up to 5% for 4-bit quantization, highlighting that QwT is particularly effective for low-bit scenarios. Additionally, after the QwT modules are integrated, the accuracy in 6-bit quantization cases aligns closely with prior state-of-the-art approaches [38, 47].
@@ Appendix B More Experimental Results
Table 11: Full results on the ResNet [17] backbone.
@@
Percentile[23] 4/4 006.1 58.3
Percentile + QwT 4/4 006.4 68.9
Percentile + QwT 4/4 006.4 69.4
@@
Percentile[23] 4/4 014.0 68.4
Percentile + QwT 4/4 016.0 74.5
Percentile + QwT 4/4 016.0 75.8
@@
Percentile[23] 6/6 019.9 76.0
Percentile + QwT 6/6 021.9 76.8
Percentile + QwT 6/6 021.9 76.8
@@
Percentile[23] 6/6 034.3 77.1
Percentile + QwT 6/6 038.6 77.2
Percentile + QwT 6/6 038.6 77.2

### v4

@@ 4 Experiments
Percentile[23] 4/4 014.0 62.3
Percentile + QwT 4/4 016.0 68.5
Percentile + QwT 4/4 016.0 72.5
@@
Percentile[23] 6/6 019.9 76.4
Percentile + QwT 6/6 021.9 76.6
Percentile + QwT 6/6 021.9 76.6
@@ 4.1 Experiments on Image Classification
Results on different backbones. Table 1 summarizes the quantization results when applying QwT across different backbone architectures. Specifically, we selected RepQ-ViT [27] and Percentile [23] as the baseline methods for the Transformer family [11, 49, 34] and ResNet [17], respectively. The results show that incorporating the QwT module consistently boosts the recognition accuracy, leading to an average increase of approximately 2.6%, and even up to 5% for 4-bit quantization, highlighting that QwT is particularly effective for low-bit scenarios. Additionally, after the QwT modules are integrated, the accuracy in 6-bit quantization cases aligns closely with prior state-of-the-art approaches [38].
@@ 5 Conclusions
Correction Note (July 2025)
In the originally published camera-ready version of this paper, an issue in the implementation of the Percentile PTQ baseline led to incorrect configuration labels in Table 1 for the ResNet-50 backbone. The results reported under “4/4” and “6/6” bit-width settings were actually obtained using “4/8” and “6/8” configurations, respectively. We have now corrected Table 1 to display the right “4/4” and “6/6” results. Furthermore, in Table 11, we have extended our experiments to include all weight and activation bit-width combinations (“4/4”, “4/8”, “6/6”, and “6/8”) for completeness. These updates do not affect any of the paper’s conclusions. We apologize for any confusion caused.
@@ Appendix B More Experimental Results
Table 11: Full results on the ResNet [17] backbone. ‘#Bits’ indicates the bit-width of weights/activations. $\dagger$’ indicates that the previous state-of-the-art results are directly sourced from the papers [47] due to the unavailability of their official code implementations.
@@
Percentile[23] 4/4 006.1 47.1
Percentile + QwT 4/4 006.4 62.3
Percentile + QwT 4/4 006.4 65.5
@@
Percentile[23] 4/4 014.0 62.3
Percentile + QwT 4/4 016.0 68.5
Percentile + QwT 4/4 016.0 72.5
@@
Percentile + QwT 4/8 016.0 74.5
Percentile + QwT 4/8 016.0 75.8
[…]
Percentile + QwT 6/6 021.9 76.6
Percentile + QwT 6/6 021.9 76.6
[…]
Percentile + QwT 6/8 021.9 76.8
Percentile + QwT 6/8 021.9 76.8
@@
Percentile + QwT 4/8 028.0 76.4
Percentile + QwT 4/8 028.0 76.7
[…]
Percentile + QwT 6/6 038.6 76.9
Percentile + QwT 6/6 038.6 76.9
[…]
Percentile + QwT 6/8 038.6 77.2
Percentile + QwT 6/8 038.6 77.2